# Reputation in the Shadow of Exit

### v4

@@ 6 Discussion
$\delta[V^{*}(\delta)L(c^{*})-\underline{V}_{0}(\delta)]\geq v(0)-v(c^{*})L(c^{*})$
and asymptotic $c^{*}$ incentive compatibility can be written as
$\delta[V^{*}(\delta)L(c^{*})-\underline{V}_{0}(\delta)]>v(0)-v(c^{*})L(c^{*})$
because dependence on $a$ is suppressed in the likelihood ratio and $|\mathcal{N}(0)|$ is a singleton so $\underline{V}_{0}(\delta)=\bar{V}_{0}(\delta)$. These are the same incentive compatibility constraints modulo the equality constraint. But because in the folk theorem the right hand side is constant in $\delta$ but the left hand side is increasing, there must exist $\delta$ large enough where the first constraint holds strictly. Hence the two statements are equivalent.
That there is a unique Markovian equilibrium when the IC constraint is satisfied then follows by noting that (1) there is a unique static Nash equilibrium (and hence a unique strategy at all histories in an MPE where $\mu(h^{t})(\omega_{R})=1$) and (2) that $\Omega(\delta)=\{0,c^{*}(\delta)\}$ contains only two elements, so any histories where $\mu(h^{t})(\omega_{R})<1$ must be clean with a fixed element. Thus the unique MPE is the one that plays $c^{*}(\delta)\in\Omega(\delta)$ at every history. ∎
@@
Fix $\varepsilon>0$ and preferences $(u,v)$. Fix some action $c$ and $a^{*}(c)$ be its lower Stackelberg best response where $v(a^{*}(c),c)>0$, which exists by the nontriviality assumption. Set $f(a^{*}(c),c)=1-\eta,f(a^{*}(c),0)=\eta$. Let $\underline{a}=a^{*}(0)\in\mathcal{N}(0)$ be such that $(\underline{a},0)$ gives the long-run player payoff $\underline{V}_{0}(\delta)$. Set moreover $f^{*}(\underline{a},0)=f^{*}(\underline{a},c)=\eta$. Note for $\eta$ sufficiently small, playing $c$ must be preferred to playing 0 since
$\frac{v(a^{*}(c),c)(1-\eta)}{1-\delta(1-\eta)}\geq\frac{v(a,0)\eta}{1-\delta\eta}$
@@
Let $c^{*}(\delta,\eta)$ be the optimal commitment posture given discount rate $\delta$ and exit function given as above (and $V_{*}(\delta,\eta)$ its associated Stackelberg payoff). In this case, Proposition 2 implies the folk theorem is equivalent to
$f(a^{*}(c^{*}(\delta,\eta)),c^{*})[v(a^{*}(c^{*}(\delta,\eta)),c^{*})+\delta V_{*}(\delta,\eta)]\geq f(a^{*},0)[v(a^{*}(0)+\delta\underline{V}_{0}(\delta)]$
@@
$\frac{f(a^{*}(c),c)}{f(a^{*},0)}[v(a^{*}(c),c)+\delta V_{*}(\delta,\eta)]\geq[v(a^{*}(0)+\delta V_{*}(\delta)]$
$\iff$ $\frac{1-\eta}{\eta}[v(a^{*}(c),c)+\delta V_{*}(\delta,\eta)]\geq[v(a^{*}(0)+\delta\underline{V}_{0}(\delta)]$
which is clearly true as $\eta\to 0$. Hence under this technology asymptotic $c^{*}$ optimality is satisfied for all $\eta$ sufficiently small.
Fix now some $\eta$ sufficiently small for the folk theorem to hold (in particular so that $V^{*}(\delta)$ can be sustained as an equilibrium) under $f$ and consider the following (non-Markovian) strategy profile at time $t=0$:
@@
    Have the short-run player’s play to $a^{*}(\mu_{0}(\omega_{c^{*}}c^{*}))$, i.e. best respond to their prior belief supposing the rational type plays 0.
@@
    Regardless of the long-run player’s action, given them a payoff of $V^{*}(\delta)$ tomorrow.
Clearly $V^{*}(\delta)$ is sustainable by the folk theorem tomorrow; and because the rational long-run player’s action today does not affect their future continuation payoff, they will play their dominant action 0. Finally, Condition (2) is exactly the condition for short-run players to myopically best reply to this strategy by the long-run player. As $\mu_{0}(\omega_{R})\to 1$, short-run players (by upper hemi-continuity) converge to an action in $\mathcal{N}(0)$. In particular, this implies $\underline{a}$ must be short-run players a (weakly) lower payoff than $\lim\limits_{\alpha\to 0}v(a^{*}(\alpha c^{*}),c^{*})$.
@@
$f(\underline{a},c)[v(\underline{a},0)+\delta V_{*}(\delta)]=\eta[v(\underline{a},0)+\delta V_{*}(\delta)].$
Choosing $\eta<\frac{\varepsilon}{v(\underline{a},0)+\delta V_{*}(\delta)}$ then implies the argument. ∎
@@
For all $\varepsilon>0$, there exists $\bar{\tau}>0$, $\bar{\delta}<1$, and $\bar{\varsigma}(\tau,\delta)$ vanishing as $\tau\to 0,\delta\to 1$, such that if $\text{var}(P_{\theta})=\tau<\bar{\tau}$, $\delta>\bar{\delta}$, and $\varsigma<\bar{\varsigma}(\tau,\delta)$, then
  1. (1)
    If $R(\mu)<0$, asymptotic $c^{*}$ incentive compatibility holds, $\underline{U}_{0}^{M}(\delta)\geq V^{*}(\delta)-\varepsilon$ and $f(a^{*}(c^{*}))>1-\varepsilon$.
  2. (2)
    If instead $R(\mu)>0$, the $\max\{\underline{U}_{0}(\delta),\bar{U}^{M}(\delta)\}<\varepsilon$.
@@
because $F_{x}(\cdot|\theta)$ converges to $\delta_{\theta}$ as $\varsigma\to 0$ (people become arbitrarily sure about the state). Since $\theta^{*}(c)$ is continuous and zeros $F_{x}\left(\frac{1}{1+c^{*}}|\theta\right)-\theta$, this implies $\theta^{*}(c)=\frac{1}{1+c^{*}}$.
@@
$\delta(V^{*}(\delta)-\underline{V}_{0}(\delta))>\kappa(c^{*})$
in the right order of limits. Some algebra implies where $\bar{P}_{\theta}(x)=1-P_{\theta}(\theta^{*}(x))$ is the probability of survival given a commitment posture. Taking $\delta$ sufficiently large and noting the inequality is strict implies that this statement is equivalent to the requirement
$\bar{P}_{\theta}(c^{*})-\bar{P}_{\theta}(0)\geq\kappa(c^{*})[1-\bar{P}_{\theta}(0)].$
@@
$\bar{P}_{\theta}(c^{*})=1-P_{\theta}(\theta^{*}(c^{*}))>1-P_{\theta}(\mu-\eta)\geq 1-\frac{\text{var}(P_{\theta})}{\eta^{2}}>1-\frac{\bar{\tau}}{\eta}.$
The middle inequality uses Chebychev’s inequality. Thus $\bar{P}_{\theta}(c^{*})\to 1$ as $\bar{\tau}\to 0$; similarly $\bar{P}_{\theta}(0)\to 0$ as $\bar{\tau}\to 0$; together this implies the incentive constraint holds. Theorem 1 and Corollary 1 imply the regime can sustain uniquely $c^{*}$ and some best response to $c^{*}$ in the unique Markovian equilibrium, as desired. Finally because $\bar{P}_{\theta}(c^{*})=f(c^{*})$, the probability of exit vanishes as $\bar{\tau}\to 1$. This implies both parts of the first statement, noting $\bar{\varsigma}$ can depend on the choice of $(\tau,\delta)$.
@@
$\frac{\bar{P}_{\theta}(c^{*})-\bar{P}_{\theta}(0)}{1-\delta\bar{P}_{\theta}(0)}>\kappa(c^{*})$
For any fixed $\delta$, note that as $\tau$ vanishes, one has that
$1-P_{\theta}(\theta^{*}(\sigma^{*}(h^{t}))\leq 1-P_{\theta}(\theta^{*}(c^{*}))<1-P_{\theta}(\mu+\eta)\leq\frac{\text{var}(P_{\theta})}{\eta^{2}}<\frac{\bar{\tau}}{\eta^{2}}$
@@
again using Chebychev’s inequality. Since $\theta^{*}(c)<\theta^{*}(0)$, we can conclude the numerator must vanish while the denominator converges to $1-\delta$. Hence this inequality will fail strictly for all sufficiently small $\tau$ for any fixed $\delta$. This implies that $c^{*}$-incentive compatibility must fail strictly, so by the converse Theorem 1 in the unique Markovian equilibrium the regime plays 0 in every period. Since $\bar{P}_{\theta}(0)$ vanishes, $\underline{U}_{0}(\delta)<\varepsilon$. Since this was the unique Markovian equilibrium, $\underline{U}^{M}(\delta)<\varepsilon$ as well. This finishes both directions of the proof. ∎

### v5

@@ 6 Discussion
$(V^{*})\mathcal{L}(c^{*})-\underline{V}_{0}\geq v(0)-v(c^{*})\mathcal{L}(c^{*})$
and $c^{*}$ incentive compatibility can be written as
$(V^{*})\mathcal{L}(c^{*})-\underline{V}_{0}>v(0)-v(c^{*})\mathcal{L}(c^{*})$
because dependence on $a$ is suppressed in the likelihood ratio and $|\mathcal{N}(0)|$ is a singleton so $\underline{V}_{0}=\overline{V}_{0}$. These are the same incentive compatibility constraints. Hence the two statements are equivalent.
That there is a unique Markov equilibrium when the IC constraint is satisfied then follows by noting that (1) there is a unique static Nash equilibrium (and hence a unique strategy at all histories in an MPE where $\mu(h^{t})(\omega_{R})=1$) and (2) that $\Omega=\{0,c^{*}\}$ contains only two elements, so any histories where $\mu(h^{t})(\omega_{R})<1$ must be clean with a fixed element. Thus the unique MPE is the one that plays $c^{*}\in\Omega$ at every history. ∎
@@
Fix $\varepsilon>0$ and preferences $(u,v)$. Fix some action $c$ and $a^{*}(c)$ be some best reply to it where $v(a^{*}(c),c)>0$, which exists by the nontriviality assumption. Set $f(a^{*}(c),c)=1-\eta,f(a^{*}(c),0)=\eta$. Let $\underline{a}=a^{*}(0)\in\mathcal{N}(0)$ be such that $(\underline{a},0)$ gives the long-run player payoff $\underline{V}_{0}$. Set moreover $f^{*}(\underline{a},0)=f^{*}(\underline{a},c)=\eta$. Note for $\eta$ sufficiently small, playing $c$ must be preferred to playing 0 since
$\frac{v(a^{*}(c),c)(1-\eta)}{1-(1-\eta)}\geq\frac{v(a,0)\eta}{1-\eta}$
@@
Let $c^{*}(\eta)$ be the optimal commitment posture given the exit function described above (and $V_{*}(\eta)$ its associated Stackelberg payoff). In this case, Proposition 1 implies the folk theorem is equivalent to
$f(a^{*}(c^{*}(\eta)),c^{*})[v(a^{*}(c^{*}(\eta)),c^{*})+V_{*}(\delta,\eta)]\geq f(a^{*},0)[v(a^{*}(0)+\underline{V}_{0}]$
@@
$\frac{f(a^{*}(c),c)}{f(a^{*},0)}[v(a^{*}(c),c)+V_{*}(\eta)]\geq[v(a^{*}(0)+\delta V_{*}]$
$\iff$ $\frac{1-\eta}{\eta}[v(a^{*}(c),c)+V_{*}(\eta)]\geq[v(a^{*}(0)+\underline{V}_{0}]$
which is clearly true as $\eta\to 0$. Hence under this survival probability function $f^{*}$, $c^{*}$ optimality is satisfied for all $\eta$ sufficiently small.
Fix now some $\eta$ sufficiently small for the folk theorem to hold (in particular so that $V^{*}$ can be sustained as an equilibrium) under $f$ and consider the following (non-Markov) strategy profile at time $t=0$:
@@
    Have the short-run player’s play to $a^{*}(\mu_{0}(\omega_{c^{*}})c^{*}))$, i.e. best respond to their prior belief supposing the rational type plays 0.
@@
    Regardless of the long-run player’s action, given them a payoff of $V^{*}$ tomorrow.
Clearly $V^{*}$ is sustainable by the folk theorem tomorrow; and because the rational long-run player’s action today does not affect their future continuation payoff, they will play their dominant action 0. Finally, Condition (2) is exactly the condition for short-run players to myopically best reply to this strategy by the long-run player. As $\mu_{0}(\omega_{R})\to 1$, short-run players (by upper hemi-continuity) converge to an action in $\mathcal{N}(0)$. In particular, this implies $\underline{a}$ must be short-run players a (weakly) lower payoff than $\lim\limits_{\alpha\to 0}v(a^{*}(\alpha c^{*}),c^{*})$.
@@
$f(\underline{a},c)[v(\underline{a},0)+V_{*}]=\eta[v(\underline{a},0)+V_{*}].$
Choosing $\eta<\frac{\varepsilon}{v(\underline{a},0)+\delta V_{*}}$ then implies the argument.
Finally, the long-run player’s payoff can be arbitrarily high. For fixed $\eta$, set $f(a^{*}(c^{*}(\eta),c^{*}(\eta)))=1-\varsigma$. Then
$V^{*}\geq(1-\varsigma)\frac{v(a^{*}(c^{*}(\eta)),c^{*}(\eta))}{\varsigma}$
which diverges as $\varsigma\to\infty$. ∎
PROOF OF PROPOSITION 4
Proof.
The first equivalence. Rank $F\succ A$ and $In\succ Out$. Then $f(a,c)$ and $v(a,c)$ are strictly decreasing for each fixed $c$, and $\mathcal{N}(\alpha\text{Fight})$ is decreasing in $\alpha$, so the game is monotone. Thus the first statement follows from Theorem 1 once we verify $c^{*}$-incentive compatibility. First, note that
$V_{*}=\frac{2(1-\eta)}{1-(1-\eta)}\text{ and }\overline{V}_{0}=\frac{\varepsilon}{1-\varepsilon}.$
When $a=Out$, $L(a,c^{*})=1$; hence the condition boils down to
$\left(\frac{2(1-\eta)}{1-(1-\eta)}-\frac{\varepsilon}{1-\varepsilon}\right)>(2-2)$
which is clearly satisfied since $1-\eta>\varepsilon$. When instead $a=In$, the condition becomes
$\left(\frac{2(1-\eta)}{1-(1-\eta)}\frac{\varphi}{\varepsilon}-\frac{\varepsilon}{1-\varepsilon}\right)>1$
Clearly the expression on the right hand side is decreasing in $\varepsilon$ and satisfied as $\varepsilon\to 0$ for fixed $\varphi$, and so in particular the intermediate value theorem furnishes the necessary $\varepsilon$ (possibly $\varepsilon=1$) for the result to hold.
The second part of the result follows by constructing the type of equilibrium in Proposition 3, taking $\varepsilon\to 0$ and noting that the long-run player will not fight in the first period because the “exit probability” is sunk. ∎
@@
because $F_{x}(\cdot|\theta)$ converges to $\delta_{\theta}$ as $\varsigma\to 0$ (people become extremely sure about the state). Since $\theta^{*}(c)$ is continuous and zeros $F_{x}\left(\frac{1}{1+c^{*}}|\theta\right)-\theta$, this implies $\theta^{*}(c)=\frac{1}{1+c^{*}}$.
@@
$V^{*}-\underline{V}_{0}>\kappa(c^{*})$
in the right order of limits. Set $\overline{P}_{\theta}(x)=1-P_{\theta}(\theta^{*}(x))$ to be the probability of survival given a commitment posture. Some algebra implies that incentive compatibility is equivalent to the requirement
$\overline{P}_{\theta}(c^{*})-\overline{P}_{\theta}(0)\geq\kappa(c^{*})[1-\overline{P}_{\theta}(0)].$
@@
$\overline{P}_{\theta}(c^{*})=1-P_{\theta}(\theta^{*}(c^{*}))>1-P_{\theta}(\mu-\eta)\geq 1-\frac{\text{var}(\theta)}{\eta^{2}}>1-\frac{\overline{\tau}}{\eta}.$
The middle inequality uses Chebychev’s inequality. Thus $\overline{P}_{\theta}(c^{*})\to 1$ as $\overline{\tau}\to 0$; similarly $\overline{P}_{\theta}(0)\to 0$ as $\overline{\tau}\to 0$; together this implies the incentive constraint holds. Theorem 1 and Corollary 1 imply the regime can sustain $c^{*}$ and some best response to $c^{*}$ in the unique Markov equilibrium, as desired. Finally because $\overline{P}_{\theta}(c^{*})=f(c^{*})$, the probability of exit vanishes as $\overline{\tau}\to 1$. This implies both parts of the first statement, noting $\overline{\varsigma}$ can depend on the choice of $\tau$.
@@
$\frac{\overline{P}_{\theta}(c^{*})-\overline{P}_{\theta}(0)}{1-\overline{P}_{\theta}(0)}>\kappa(c^{*})$
As $\bar{\tau}$ vanishes, one has that
$1-P_{\theta}(\theta^{*}(\sigma^{*}(h^{t}))\leq 1-P_{\theta}(\theta^{*}(c^{*}))<1-P_{\theta}(\mu+\eta)\leq\frac{\text{var}(\theta)}{\eta^{2}}<\frac{\overline{\tau}}{\eta^{2}}$
again using Chebychev’s inequality. Since $\theta^{*}(c)<\theta^{*}(0)$, we conclude that
$\frac{\overline{P}_{\theta}(c^{*})-\overline{P}_{\theta}(0)}{1-\overline{P}_{\theta}(0)}\leq\frac{\frac{\bar{\tau}}{\eta^{2}}-\bar{P}_{\theta}(0)}{1-\bar{P}_{\theta}(0)}.$
@@
Hence this term is bounded from above by $\kappa(c^{*})$ for all sufficiently small $\tau$. Thus, $c^{*}$-incentive compatibility must fail strictly, so by the converse Theorem 1 in the unique Markov equilibrium the regime plays 0 in every period. Since $\overline{P}_{\theta}(0)$ vanishes, $\underline{U}_{0}<\varepsilon$. As this was the unique Markov equilibrium, $\underline{U}^{M}<\varepsilon$ as well. This finishes both directions of the proof. ∎
SUPPLEMENTAL APPENDIX
This appendix considers the discounted repeated game at some discount rate $\delta$. Section B.1 discusses some problems with normalized repeated games, interpreted as the average payoffs across periods. Section B.2 derives a novel comparative statics result on the set of commitment postures as the discount rate changes, and interprets it in the context of applications. Section B.3 sets the main result in the language of the discounted game.
B.1. Non-normalized payoffs
Intuitively, without exit, the stream of normalized discounted payoffs converges to the average payoff as $\delta\to 1$. However, when the game has endogenous exit, so long as exit occurs with high enough probability, the average stream of payoffs will be zero for “most” periods (i.e. if the expected exit time is finite), so normalized payoffs will vanish as players get extremely patient. The lemma formalizes this observation.
Lemma B.1.
$\lim\limits_{\delta\to 1}(1-\delta)\sum_{t=0}^{\infty}v_{t}\prod_{s=0}^{t}f_{s}=0.$
for any feasible payoffs $v_{t}$.
Proof.
Let $\{\overline{f}_{\tilde{t}}\}\subset\{f_{t}\}$ be a subsequence such that $\lim\limits_{\tilde{t}\to\infty}\overline{f}_{\tilde{t}}=\overline{f}<1-\eta$, for some $\eta>0$ fixed; moreover, by truncating finitely many terms at most, suppose $\overline{f}_{\tilde{t}}<1-\eta$ for all $s\in\mathbb{N}$. Now for any $T\in\mathbb{N}$, define the number of times that $f_{t}<1-\eta$ to be $s(T)=\#\{s:f_{s}<1-\eta\}$. By assumption of $\eta$, we know $\lim\limits_{T\to\infty}s(T)=\infty$. We now have that $\prod_{s=0}^{\infty}f_{t}<(1-\eta)^{s(T)}$. Letting $\overline{v}$ be chosen so that $|v_{t}|\leq\overline{v}$ for all $t$, we have
$0$ $\leq\lim\limits_{\delta\to 1}\left|(1-\delta)\sum_{t=0}^{\infty}v_{t}\prod_{s=0}^{\infty}f_{s}\right|\leq\lim\limits_{\delta\to 1}\left|(1-\delta)\sum_{t=0}^{\infty}\overline{v}(1-\eta)^{s(t)}\right|$
$\leq\lim\limits_{\delta\to 1}\left|(1-\delta)\sum_{t=0}^{T}\overline{v}(1-\eta)^{s(t)}\right|+\left|(1-\delta)(1-\eta)^{s(T)}\sum_{t=T+1}^{\infty}\overline{v}\right|$
$\leq(1-\eta)^{s(T)}\lim\limits_{\delta\to 1}(1-\delta)\sum_{t=T+1}^{\infty}\overline{v}\leq(1-\eta)^{s(T)}\overline{v}.$
Lemma B.1 implies that it is not fruitful to compare discounted average payoffs in the patient limit, as they will often be zero even when payoffs are meaningfully different.
This choice, however, leads itself to a novel technical challenge: without normalizing payoffs by the discount rate, deviations today may remain first order (in particular, they are not order $(1-\delta)$ compared to an order $\delta$ variation in the continuation payoff, and are instead order $1$ versus order $\delta$ variation in the continuation payoff). This makes it much harder to discipline the long-run player from deviating to play non-Stackelberg actions, since deviations from potential equilibrium profiles will always remain tempting. Theorem 1 can thus be interpreted as providing a simple economic condition under which the long-run player can successfully establish a (Markov) reputation, despite this hurdle.
B.2. Commitment Postures
How does the discount rate affect the long-run player’s Stackelberg payoff? Define
$c^{*}=\operatorname*{\arg\!\max}_{c\in C}\min_{a\in\mathcal{N}(c)}\sum_{t=0}^{\infty}f(a,c)^{t+1}v(a,c)$
to be the set of (lower) Stackelberg actions for the long-run player when they discount the future at rate $\delta$. Note the long-run player trades off between a higher survival probability at the (potential) cost of a lower stage-game payoff in each period. My next result characterizes exactly how the long-run player resolves this tradeoff as they get more patient: regardless of the speed at which $v(a,c)$ might decline in $c$, they will always favor more aggressive actions that increase their survival probability as they get more patient. To do so, I make one more assumption on the survival probability.
Assumption 3.
$f^{*}(c)=\min_{a\in\mathcal{N}(c)}f(a,c)$ is nondecreasing in $c$.
Proposition B.1 (Patience Effect).
Suppose Assumption 3 is satisfied. $c^{*}$ is nonempty, upper hemicontinuous, and increasing (in the strong set order) in $\delta$.
Proof.
We first claim $f^{*}(c)$ is continuous, since for any sequence $c_{n}\to c$,
$\lim\limits_{n\to\infty}f^{*}(c_{n})=\lim\limits_{n\to\infty}\frac{v(a^{*}(c_{n}),c_{n})f(a^{*}(c_{n}),c_{n})}{1-f(a^{*}(c_{n}),c_{n})}=\frac{v(a^{*}(c),c)f(a^{*}(c),c)}{1-f(a^{*}(c),c)}.$
The first inequality is by definition, and the second by upper hemi-continuity of the Nash correspondence. Given continuity of $f^{*}(c)$, Berge’s theorem of the maximum applied to the optimizing set $c^{*}=\max_{c\in C}f^{*}(c)$ implies $c^{*}$ exists and is upper hemi-continuous. Monotonicity now relies on the following lemma [^18] Such a lemma appears elementary, but to the best of my knowledge there is not a prior reference. Note the lemma establishes a general comparative statics result for discounted streams of payoffs that cross zero once from below..
Lemma B.2.
Proof.
Note
$\sum_{t=0}^{\infty}\delta_{1}^{t}b_{t}=\sum_{t=0}^{\infty}\delta_{1}^{t-T}b_{t}\geq 0\iff\sum_{t=0}^{\infty}\delta_{1}^{t-T}b_{t}$
Fix any $\delta$ and $c^{\prime}>c$; by Assumption 3, it must be that for all $(c^{\prime},c)$, $f^{*}(c^{\prime}),f^{*}(c)>0$. Consider the difference
$\sum_{t=0}^{\infty}[v(a^{*}(c^{\prime}),c^{\prime})f^{*}(c^{\prime})^{t+1}-v(a^{*}(c),c)f^{*}(c)^{t+1}]$
then
The left hand side is strictly positive, while the right hand side is positive, in $(0,1)$ (by Assumption 3), and strictly decreasing in $t$. Thus, there exists a unique $T$ (possibly 0) at which the inequality holds for all $t\geq T$. Lemma B.2 then implies, for any $c^{\prime}>c$, $\sum_{t=0}^{\infty}b_{t}$ is increasing in $\delta$, and hence the objective function $\sum_{t=0}^{\infty}v(a^{*}(c),c)$ is single-crossing in $(\delta,c)$. The result then follows by the Milgrom and Shannon (1994) comparative statics theorem. ∎
Proposition B.1 applies regardless of the speed or curvature of $v$ (in particular, it holds even if $v$ declines very fast relative to $c$), though this comes at the cost of a weak comparative static (every selection from $c^{*}$ is nondecreasing in $c$). Moreover, it has economic implications in the context of the applications—a regime seeking to fend off a coordinated attack will invest more in survival when they are more patient; longer-lived firms care more about ensures entrants are deterred; patient motorcyclists have a greater incentive to take “separating” actions that prove they are not commitment types, in order to deter exit.
B.3. The Main Result
That $c^{*}$ varies with the discount rate produces a new order of limits challenge in trying to prove some version of a commitment payoff theorem in the discounted repeated game. In particular, since the long-run player’s Stackelberg action depends on the discount rate, it is not clear that there is a finite set $\Omega$ such that $\Omega\cap c^{*}$ is nonempty for all $\delta$ close to $1$. Consequently, for any fixed set of commitment types, it may be infeasible to approximate a patient committed sender.
To obviate this problem, I index and simplify the set of commitment postures, $\Omega=\{\omega_{R},\omega_{c^{*}}\}$ (where I abuse notation and use $c^{*}$ to refer to a single selection from the set of commitment types). Moreover, as before, I set $\underline{U}$ and $\overline{U}$ as the infimum and supremum of the long-run player’s best and worst sequential equilibrium payoff, though now I implicitly fix the set of commitment types $\Omega=\Omega$, though I still suppress dependence on the specific sequence of prior beliefs $\mu_{0}\in\Delta(\Omega)$ (which itself my also change in $\delta$). Give this notation, the limit infimum (supremum) $\liminf_{\delta\to 1}\underline{U}$ (resp. $\limsup_{\delta\to 1}$) should be interpreted as the limit infimum (supremum) of the sequence of lower equilibrium payoffs in the repeated game with discount rate $\delta$, commitment types $\Omega$, and some fixed sequence of prior beliefs $\mu_{0}\in\Delta(\Omega)^{o}$. This interpretation implies that statements of the form $\limsup_{\delta\to 1}|V_{*}-\underline{U}|=0$ can be read as “The long-run player attains almost their Stackelberg payoff so long as there is positive probability players believe they are committed to their Stackelberg action at their level of patience”.
Under this interpretation, the following definition and result are immediate applications of the undiscounted game, by continuity of the payoff parameters in the discount rate.
Definition B.1.
Say the long-run player asymptotically satisfies (weak) $c^{*}$ incentive compatibility if there exists $\overline{\delta}<1$ such that for all $\delta>\overline{\delta}$ and each $a\in\mathcal{N}(\alpha c)$ for $\alpha\in[0,1]$,
$\delta\left(V_{*}(\delta)L(a,c^{*}(\delta))-\overline{V}_{0}(\delta)\right)>(\geq)\left(v(a,0)-v(a,c^{*}(\delta))L(a,c^{*}(\delta))\right)$
Theorem 3.
$\limsup_{\delta\to 1}|\underline{U}^{M}(\delta)-V_{*}(\delta)|=0.$
Conversely, if $(f,v,u)$ is monotone and the long-run player does not asymptotically satisfy weak $c^{*}$ incentive compatibility, then for all $\delta<1$,